# A relativistic treatment of accretion disk torques on extreme mass ratio inspirals around spinning black holes

### v1

@@ I Introduction
Extreme mass-ratio inspirals (EMRIs), where a small compact object (SCO) [of mass in the range $10M_{\odot}$$100M_{\odot}$] spirals into a supermassive black hole (SMBH) [of mass $10^{5}M_{\odot}$$10^{7}M_{\odot}$], emit gravitational waves in the sensitivity band of the Laser Interferometer Space Antenna (LISA). EMRIs can accumulate $\gtrsim\mathcal{O}(10^{5})$ wave-cycles in the LISA band, and thus, they provide detailed information about the spacetime in the close vicinity of the SMBH. The relativistic nature of EMRI orbits and the proximity of the SCO to the SMBH should allow one to probe the mass and spin of the SMBH, conduct tests of general relativity, and potentially probe the environment around the SMBH Seoane et al. (2023); Colpi et al. (2024); Babak et al. (2017); Kocsis et al. (2011); Yunes et al. (2011a); Speri et al. (2023); Duque et al. (2025).
@@
Accurate inference of the EMRI system parameters from a future gravitational-wave signal requires waveform models that are highly accurate because of the large number of cycles that accumulate in the LISA band Colpi et al. (2024). These high-accuracy requirements are being tackled successfully in vacuum (i.e. in the absence of any external matter source around the SMBH) thanks to advances in self-force and black hole perturbation theory Afshordi et al. (2023); BHP . With the modeling of EMRIs in vacuum reaching a mature stage, it is now time to assess the impact of environmental effects on the EMRI motion.
@@
Environmental effects have long been expected to be a small perturbation on EMRIs. This is because EMRIs are expected to exist at Gpc distances from Earth, and thus, their gravitational waves are naturally weak. In order to make them detectable, the SCO must be close enough to the SMBH, so that the SCO’s velocity is a good fraction of the speed of light and the gravitational-wave amplitude is not suppressed. The matter-energy density in the vicinity of SMBHs is expected to be small relative to the matter-energy of the SMBH, and thus, any (non-vacuum) environmental effects may be thought to be negligible. However, the large number of gravitational-wave cycles EMRIs will deposit in the LISA band will allow us to measure small perturbations from the environment, if the dephasings caused are on the order of a few radians or higher Yunes et al. (2011a); Kocsis et al. (2011).
@@
A large body of work has now qualitatively assessed how different environmental effects could be potentially studied with EMRI signals. A non-exhaustive list of astrophysical environmental effects include the interaction between the SCO and the accretion disk of an SMBH Narayan (2000); Levin (2007); Kocsis et al. (2011); Yunes et al. (2011a); Speri et al. (2023); Duque et al. (2025); Derdzinski et al. (2019, 2021), mass accretion on to the SMBH and the SCO Kocsis et al. (2011); Yunes et al. (2011a), viscous drag on the SCO due to gas surrounding the SCO Barausse and Rezzolla (2008), disk-SCO collisions Franchini et al. (2023), and perturbations due to the presence of a third body Yunes et al. (2011b); Naoz et al. (2013); Gupta et al. (2021); Silva and Hirata (2022); Silva et al. (2025). In addition, there could be scenarios where the SCO interact with bosonic clouds and dark matter particles surrounding a SMBH Baumann et al. (2022); Maselli et al. (2020); Zhang and Yang (2020); Arvanitaki and Dubovsky (2011); Vicente and Cardoso (2022); Dyson et al. (2025), or where EMRIs are part of a three-body system Yunes et al. (2011b); Santos et al. (2025); Yin et al. (2025); Cocco et al. (2025); Camilloni et al. (2024).
@@
Several studies have assessed the potential for EMRIs to probe disk-SCO interactions Yunes et al. (2011a); Kocsis et al. (2011); Sberna et al. (2022); Speri et al. (2023); Copparoni et al. (2025); Duque et al. (2025); Derdzinski et al. (2019, 2021); however, except for Hirata (2011a, b), all of these studies have used the disk-satellite formula derived in the proto-planetary literature to model the disk-SCO interaction. In addition, the relativistic model introduced in Hirata (2011a, b) (and studied more recently in Franscisco Duque and Vicente (2025, in prep.)) is a purely numerical approach that relies on the integration of the Teukolsky equation. In an effort to build simple relativistic models of disk-SCO interaction close to a SMBH, we introduced an analytical model in Hegade K. R. et al. (2025) (Paper I, hereafter), using tools from self-force and Hamiltonian perturbation theory. The methods introduced in Paper I model the interaction between the disk and the SCO as a three-body problem, where a disk particle close to the orbit of the SCO exchanges energy and angular momentum with the SCO at Lindblad resonances. Such simplified methods were used to study Newtonian proto-planetary disks in the past Goldreich and Tremaine (1980, 1981); Ogilvie (2007); Ward (1986, 1997), and our approach extended these methods to a Schwarzschild (non-rotating black hole) background setting.
@@
These effects have been confirmed by an independent work Franscisco Duque and Vicente (2025, in prep.) that follows the approach of Hirata (2011a, b).
@@ II Qualitative discussion of the impact of spin on torque reversal
Consider a SCO of mass $m$ on an equatorial, circular orbit around a SMBH of mass $M$ at radius $r=p^{\prime}M$ [^1] A prime denotes the orbital elements of the SCO, unless otherwise stated. in Boyer-Lindquist coordinates that interacts with a thin, equatorial disk of surface density $\Sigma(r)$. We assume that the angular momentum of the disk and the CO point in the same direction, but we let the angular momentum of the SMBH be arbitrary. The dominant source of torque for the disk-SCO configuration described above arises from the exchange of energy and angular momentum at Lindblad resonances Goldreich and Tremaine (1980). The location of Lindblad resonances $r_{\mathrm{Lr}}=p_{\mathrm{Lr}}M$ in Newtonian gravity are given by the roots of Goldreich and Tremaine (1980); Ward (1986)
@@ III.1.1 Step 1: SCO-Disk Interaction
where $h_{\mu\nu}^{\mathrm{S,SCO}}(z;z^{\prime})$ and $h_{\mu\nu}^{\mathrm{R,SCO}}(z;z^{\prime})$ are the singular and the regular field contributions from the SCO. In the second line of the above equation, we have assumed that the Newtonian-like singular potential of the SCO $h_{\mu\nu}^{\mathrm{S,SCO}}(z;z^{\prime})$ is the dominant contribution to the interaction, and we ignored the regular field of the SCO by assuming that the interaction is quasi-adiabatic. Using analytic expressions for the singular field Pound and Miller (2014); Heffernan et al. (2012) one can then evaluate the above equation (see Paper I for further details).
@@ III.2 Lindblad resonances
for a circular orbit. We note that the value of $\left<\dot{E}^{\prime}\right>_{\mathrm{Lr,k}}^{(0)}$ can also be obtained by a calculation in the local Fermi frame, see Sec. IV C of Paper I and this is a consistency check on our calculatio.
@@ IV.3 Comparison to gravitational wave emission
In Fig. 5, we plot this ratio for a SMBH with mass $M=10^{6}M_{\odot}$ surrounded by an accretion disk with $\Sigma_{h}=1$ and $\Sigma_{0}=10^{3}$, $\Sigma_{p}=-3/2$ ($\alpha$ disk-like values) on the left panel and $\Sigma_{0}=10^{5}$, $\Sigma_{p}=-3/5$ ($\beta$ disk-like values) on the right panel. In both panels, we choose a few different values of SMBH spin and accretion disk aspect ratio. For the gravitational wave flux, we use the analytical expressions from Fujita (2015), as coded in Black Hole Perturbation Toolkit BHP . As we see from the figure, the value of the energy flux due to the disk-SCO interaction is always smaller than the loss of energy to gravitational-wave emission. Therefore, the torque reversal cannot cause a floating orbit, unless $\Sigma_{0}$ is extremely large. We also see that, for the $\chi=0.9$ case, the torque is smaller than the $\chi=0$ and $\chi=-0.9$ values; this is because the ISCO is very close to the horizon, and therefore, the energy loss due to gravitational wave emission is always significantly larger. For a retrograde SCO orbit and a retrograde disk ($\chi=-0.9$), we see that the torque is only 1-3 orders of magnitude smaller than the gravitational-wave contribution, because the ISCO approaches $9M$ as $\chi\to-1$. This suggests that for retrograde orbits, the disk-SCO interaction will have a significantly larger impact on the waveform. Finally, although the ratio is very small, the large amount of cycles spent by an EMRI in the LISA band could help us disentangle the disk-SCO effect from the vacuum EMRI signal. We refer the reader to Paper I for a heuristic dephasing estimate in a Schwarzschild background spacetime. For Bayesian and Fisher estimates using the Newtonian torque formula, see Yunes et al. (2011a); Kocsis et al. (2011); Speri et al. (2023).
@@ V Conclusions
In this paper, we extended the results of Paper I by analyzing the relativistic treatment of disk-SCO interactions to a Kerr background spacetime. For nearly circular and equatorial orbits in a Kerr spacetime, we provided a detailed analytic exploration of the effects of the spin on the torque due to Lindblad resonances. Our two key results are the following. First, we provided analytical expressions for the power and torque induced on the SCO from the disk-SCO interaction in an EMRI [Eqs. (18) and (20)]. Second, we investigated the impact of the SMBH spin and the accretion disk surface density gradient on the torque reversal location [ Figs. 2 and 3]. We have found that the spin of the Kerr background tends to enhance the relativistic disk-SCO interaction, and this must be taken into account in future EMRI modeling of environmental effects. Some of these results have been confirmed by an independent work Franscisco Duque and Vicente (2025, in prep.), albeit numerically.

### v3

@@ I Introduction
Extreme mass-ratio inspirals (EMRIs), where a small compact object (SCO) [of mass in the range $10M_{\odot}$$100M_{\odot}$] spirals into a supermassive black hole (SMBH) [of mass $10^{5}M_{\odot}$$10^{7}M_{\odot}$], emit gravitational waves in the sensitivity band of the Laser Interferometer Space Antenna (LISA). EMRIs can accumulate $\gtrsim\mathcal{O}(10^{5})$ wave-cycles in the LISA band, and thus, they provide detailed information about the spacetime in the close vicinity of the SMBH. The relativistic nature of EMRI orbits and the proximity of the SCO to the SMBH should allow one to probe the mass and spin of the SMBH, conduct tests of general relativity, and potentially probe the environment around the SMBH Seoane and others (2023); Colpi and others (2024); Babak et al. (2017); Kocsis et al. (2011); Yunes et al. (2011b); Speri et al. (2023); Duque et al. (2025a).
@@
Accurate inference of the EMRI system parameters from a future gravitational-wave signal requires waveform models that are highly accurate because of the large number of cycles that accumulate in the LISA band Colpi and others (2024). These high-accuracy requirements are being tackled successfully in vacuum (i.e. in the absence of any external matter source around the SMBH) thanks to advances in self-force and black hole perturbation theory N. Afshordi et al. (2023); 10. With the modeling of EMRIs in vacuum reaching a mature stage, it is now time to assess the impact of environmental effects on the EMRI motion.
@@
Environmental effects have long been expected to be a small perturbation on EMRIs. This is because EMRIs are expected to exist at Gpc distances from Earth, and thus, their gravitational waves are naturally weak. In order to make them detectable, the SCO must be close enough to the SMBH, so that the SCO’s velocity is a good fraction of the speed of light and the gravitational-wave amplitude is not suppressed. The matter-energy density in the vicinity of SMBHs is expected to be small relative to the matter-energy of the SMBH, and thus, any (non-vacuum) environmental effects may be thought to be negligible. However, the large number of gravitational-wave cycles EMRIs will deposit in the LISA band will allow us to measure small perturbations from the environment, if the dephasings caused are on the order of a few radians or higher Yunes et al. (2011b); Kocsis et al. (2011).
@@
A large body of work has now qualitatively assessed how different environmental effects could be potentially studied with EMRI signals. A non-exhaustive list of astrophysical environmental effects include the interaction between the SCO and the accretion disk of an SMBH Narayan (2000); Levin (2007); Kocsis et al. (2011); Yunes et al. (2011b); Speri et al. (2023); Duque et al. (2025a); Derdzinski et al. (2019, 2021), mass accretion on to the SMBH and the SCO Kocsis et al. (2011); Yunes et al. (2011b), viscous drag on the SCO due to gas surrounding the SCO Barausse and Rezzolla (2008), disk-SCO collisions Franchini et al. (2023), and perturbations due to the presence of a third body Yunes et al. (2011a); Naoz et al. (2013); Gupta et al. (2021); Silva and Hirata (2022); Silva et al. (2025). In addition, there could be scenarios where the SCO interact with bosonic clouds and dark matter particles surrounding a SMBH Baumann et al. (2022); Maselli et al. (2020); Zhang and Yang (2020); Arvanitaki and Dubovsky (2011); Vicente and Cardoso (2022); Dyson et al. (2025), or where EMRIs are part of a three-body system Yunes et al. (2011a); Santos et al. (2025); Yin et al. (2025); Cocco et al. (2025); Camilloni et al. (2024).
@@
Several studies have assessed the potential for EMRIs to probe disk-SCO interactions Yunes et al. (2011b); Kocsis et al. (2011); Sberna and others (2022); Speri et al. (2023); Copparoni et al. (2025); Duque et al. (2025a); Derdzinski et al. (2019, 2021); however, except for Hirata (2011a, b), all of these studies have used the disk-satellite formula derived in the proto-planetary literature to model the disk-SCO interaction. In addition, the relativistic model introduced in Hirata (2011a, b) (and studied more recently in Duque et al. (2025b)) is a purely numerical approach that relies on the integration of the Teukolsky equation. In an effort to build simple relativistic models of disk-SCO interaction close to a SMBH, we introduced an analytical model in Hegade K. R. et al. (2025) (Paper I, hereafter), using tools from self-force and Hamiltonian perturbation theory. The methods introduced in Paper I model the interaction between the disk and the SCO as a three-body problem, where a disk particle close to the orbit of the SCO exchanges energy and angular momentum with the SCO at Lindblad resonances [^1] Lindblad resonances occur when the epicylic frequency of the SCO orbit is an integer multiple of the forcing frequency of the density waves raised in the accretion disk. Such resonances are the dominant source of angular and momentum exchange in disk-SCO interactions, see Binney and Tremaine (2008); Goldreich and Tremaine (1980) for a detailed discussion.. Such simplified methods were used to study Newtonian proto-planetary disks in the past Goldreich and Tremaine (1980, 1981); Ogilvie (2007); Ward (1986, 1997), and our approach extended these methods to a Schwarzschild (non-rotating black hole) background setting.
@@
These effects have been confirmed by an independent work Duque et al. (2025b) that follows the approach of Hirata (2011a, b).
@@ II Qualitative discussion of the impact of spin on torque reversal
Consider a SCO of mass $m$ on an equatorial, circular orbit around a SMBH of mass $M$ at radius $r=p^{\prime}M$ [^2] A prime denotes the orbital elements of the SCO, unless otherwise stated. in Boyer-Lindquist coordinates that interacts with a thin, equatorial disk of surface density $\Sigma(r)$. We assume that the angular momentum of the disk and the SCO point in the same direction, but we let the angular momentum of the SMBH be arbitrary. The dominant source of torque for the disk-SCO configuration described above arises from the exchange of energy and angular momentum at Lindblad resonances Goldreich and Tremaine (1980). The location of Lindblad resonances $r_{\mathrm{Lr}}=p_{\mathrm{Lr}}M$ in Newtonian gravity are given by the roots of Goldreich and Tremaine (1980); Ward (1986)
@@ III.1.1 Step 1: SCO-Disk Interaction
where $h_{\mu\nu}^{\mathrm{S,SCO}}(z;z^{\prime})$ and $h_{\mu\nu}^{\mathrm{R,SCO}}(z;z^{\prime})$ are the singular and the regular field contributions from the SCO. In the second line of the above equation, we have assumed that the Newtonian-like singular potential of the SCO $h_{\mu\nu}^{\mathrm{S,SCO}}(z;z^{\prime})$ is the dominant contribution to the interaction, and we ignored the regular field of the SCO by assuming that the resonant interaction is quasi-adiabatic. Moreover, the regular field of the SCO will be a very high post-Newtonian contribution to the dynamics, which can be incorporated in our analysis by using numerical self-force calculations 10. With this assumption, we can use analytic expressions for the singular field Pound and Miller (2014); Heffernan et al. (2012) to evaluate the above equation (see Paper I for further details).
@@ III.2 Lindblad resonances
for a circular orbit. We note that the value of $\left<\dot{E}^{\prime}\right>_{\mathrm{Lr,k}}^{(0)}$ can also be obtained by a calculation in the local Fermi frame, see Sec. IV C of Paper I and this is a consistency check on our calculation.
@@ IV.3 Comparison to gravitational wave emission
In Fig. 5, we plot this ratio for a SMBH with mass $M=10^{6}M_{\odot}$ surrounded by an accretion disk with $\Sigma_{h}=1$ and $\Sigma_{0}=10^{3}$, $\Sigma_{p}=-3/2$ ($\alpha$ disk-like values) on the left panel and $\Sigma_{0}=10^{5}$, $\Sigma_{p}=-3/5$ ($\beta$ disk-like values) on the right panel. In both panels, we choose a few different values of SMBH spin and accretion disk aspect ratio. For the gravitational wave flux, we use the analytical expressions from Fujita (2015), as coded in Black Hole Perturbation Toolkit 10. As we see from the figure, the value of the energy flux due to the disk-SCO interaction is always smaller than the loss of energy to gravitational-wave emission. Therefore, the torque reversal cannot cause a floating orbit, unless $\Sigma_{0}$ is extremely large. We also see that, for the $\chi=0.9$ case, the torque is smaller than the $\chi=0$ and $\chi=-0.9$ values; this is because the ISCO is very close to the horizon, and therefore, the energy loss due to gravitational wave emission is always significantly larger. For a retrograde SCO orbit and a retrograde disk ($\chi=-0.9$), we see that the torque is only 1-3 orders of magnitude smaller than the gravitational-wave contribution, because the ISCO approaches $9M$ as $\chi\to-1$. This suggests that for retrograde orbits, the disk-SCO interaction will have a significantly larger impact on the waveform. Finally, although the ratio is very small, the large amount of cycles spent by an EMRI in the LISA band could help us disentangle the disk-SCO effect from the vacuum EMRI signal. We refer the reader to Paper I for a heuristic dephasing estimate in a Schwarzschild background spacetime. For Bayesian and Fisher estimates using the Newtonian torque formula, see Yunes et al. (2011b); Kocsis et al. (2011); Speri et al. (2023).
@@ V Conclusions
In this paper, we extended the results of Paper I by analyzing the relativistic treatment of disk-SCO interactions to a Kerr background spacetime. For nearly circular and equatorial orbits in a Kerr spacetime, we provided a detailed analytic exploration of the effects of the spin on the torque due to Lindblad resonances. Our two key results are the following. First, we provided analytical expressions for the power and torque induced on the SCO from the disk-SCO interaction in an EMRI [Eqs. (18) and (20)]. Second, we investigated the impact of the SMBH spin and the accretion disk surface density gradient on the torque reversal location [ Figs. 2 and 3]. We have found that the spin of the Kerr background tends to enhance the relativistic disk-SCO interaction, and this must be taken into account in future EMRI modeling of environmental effects. Some of these results have been confirmed by an independent work Duque et al. (2025b), albeit numerically.